# QUANTIZED DECODER IN LEARNED IMAGE COMPRESSION FOR DETERMINISTIC RECONSTRUCTION

### 18 Dec 2023 v1

@@ 1 INTRODUCTION
Our aim is to achieve bit-exact reconstructed images at decoder when running on different platforms. For this it is sufficient to only quantize the neural network components at the decoder. For many of the use cases of image coding, the decoder of the image codec is fixed, while the encoder design is flexible. Therefore, we quantize only the decoder in order to provide bit exact reconstruction and keep the encoder side free for complex hardware implementation. We reuse the idea in [16] for post-training quantization of weights and activations without overflow and extend it to all decoder subnetworks of the base operation point of JPEG AI VM (verification model) [17] as shown in Fig. 1. We show that by quantizing only decoder networks by reusing the method to eliminate the overflow of the 32-bit accumulator for 16-bit weights and 16-bit activations, we could achieve the deterministic decoded reconstructed images with a minor loss. Also, we provide more experiments to study the 8-bit weights with 16-bit activations, and also to compare the effect of choosing the optimal clipping range instead of clipping to the maximal range for both cases with 8-bit and 16-bit weights. In Section 2, we discuss the previous related work and compare to the related work. The overall description of the learned image coding architecture is given in Section 3.1, and its quantized components are explained in Section 3.2. In Section 3.3, the method of quantization is described. The experimental results are provided in Section 4 and conclusions are given in Section 5.
Fig. 1: High-level architecture diagram of the JPEG AI base operating point.
@@ 2 RELATED WORK
In [14, 15, 16] we see the quantization of weights and activations in entropy part as a solution to the visible cross platform decoding error, additionally in [14] the weights of the main part are quantized. In this paper we extend the post training quantization method of 16-bit weights and activations with no-overflow of 32-bit accumulator in [16]. While [16] quantizes the entropy part only, this work applies quantization to the whole decoder. We target a zero deviation and achieve exact reconstructed images at cross platform tests while keeping the performance loss compared to the floating point model minimal. Although the quantization of only the entropy part as in [16] eliminates the broken images error at the decoder, there still exists a tiny deviation of the quality of the reconstructed images due to the remaining unquantized layers of the decoder, which is not visible visually but can be measured numerically. This is not a critical problem for the image compression task, but for some error-critical applications such as medical image processing, or the video compression task, where the reconstructed image typically becomes a reference for the prediction of the other frames, it is desired to have no deviation. Also we show the results of quantizing only the hyper decoder networks. This allows exact decoding of $\hat{y}$ and can be interesting for computer vision tasks which can use decoded latent tensors $\hat{y}$ directly as input, but do not need the reconstructed image.
In [10, 11, 12, 15], both weights and activations of the whole codec are quantized. In [13] only weights of the whole codec are quantized, whereas in [14] weights of the whole codec and only activations in the entropy part are quantized. We quantize not the whole codec, but only the decoder components. It is not necessary to quantize all layers in the encoder to achieve the deterministic reconstructed image. This way we allow more complexity at the calculations but the problem of the bit-exact reconstruction is still simple to solve by a post-training quantization method. It is worth discussing that with this solution, there can be a small variation at the encoded images on different devices. For common use cases such as streaming where encoding is done once on a device, and decoding is done on many devices, we can ensure that the decoder calculations on any device result in exactly the same as with the encoder calculations.
@@ 3.1 Learned image codec
As a baseline image compression codec, we use the base operating point of JPEG AI VM software version 3.4.0 [17]. The JPEG AI VM has two different operating points, high and base, with a different trade-off between complexity and performance. The base operating point architecture consists of an entropy model without context and it has simplified layers and activation functions compared to the high operating point. The architecture of the base operating point for one colour component is shown in Fig. 1. This codec processes colour components separately. Each subnetwork consists of two separate components of subnetworks, one processing the luma channel (Y) and one processing the two chroma channels (U, V). The decoder consists of three main subnetworks, a hyper scale decoder, $h_{\sigma}$, a hyper decoder, $h_{\mu}$, and a main decoder, $g_{s}$; the subnetworks are highlighted in green in Fig. 1. There are four models which are trained with different rate-distortion (RD) loss and overall they produce five different quality points during inference. G-unit and IG-unit are the trained vectors for gain and inverse gain unit, which provide the variable rate functionality as in [18]. The JPEG AI VM software has additional tools to improve the compression quality but all these additional tools are turned off during our implementation, so the coding pipeline is reduced to the neural network layers which are shown in Fig. 2. The explicit block diagram of the light-weighted residual activation unit (ResAU) and the light-weighted residual block (ResBlock) are shown on the right-hand-side of Fig. 2.
@@ 3.2 Quantized part of the codec
We use the floating point trained models as the anchor, and apply the quantization to the decoder subnetworks. The layers which are highlighted in green in Fig. 2 are the convolutional layers to be quantized. The hyper scale decoder is the subnetwork which is calculating the entropy parameters and is shared between the encoder and decoder. From the previous work, [15], [14] and [16], we know that it is necessary to quantize the entropy part of the subnetwork in order to eliminate catastrophic decoded images. In this paper, we extend the partial quantization of the codec to all the elements of the decoder, with the aim of deterministic calculation of the reconstructed image $\hat{x}$ at the encoder and decoder side. We replace the LeakyReLU activations in the hyper decoder $h_{\mu}$ with an alternative implementation using bit shift. The subnetworks which only exist at the encoder are the main encoder $g_{a}$ and the hyper encoder $h_{a}$; these subnetworks are kept as floating point in our study.
@@ 3.3 Methodology for quantization of decoder subnetworks weights and activations
In [16] weights and activations are both 16-bit. There is one difference of the weights quantization method from [16] if we apply 8-bit quantization to weights. Instead of directly scaling the weights with the calculated factor that eliminates the overflow $2^{-k_{j}}$ as in [16], we use this as an upper bound and search the power of two of scaling factor that results in the minimal quantization error for the weights in $[0,k_{j}]$. We observe that this is necessary to minimize the clipping error. Because we clip the 8-bit weights to the interval $[-2^{7},2^{7}-1]$ instead of $[-2^{15},2^{15}-1]$ as we do for the 16-bit weights case, scaling with the maximum available scaling factor is not ideal for the smaller clipping range.
@@ 4 EXPERIMENTAL RESULTS
The trained model of the JPEG AI VM base operation point is the anchor of the experiments. Then, we apply quantization to the weights and activations to the subnetworks of this model. We use the training, validation and test data which are provided by the JPEG AI common test conditions [20] [21]. The calibration data set we use at the quantization search is a 16-image subset of the validation set. For the results evaluation, we run the inference on the JPEG AI test set consisting of 50 images. We measure the average BD-rate deviation from the floating point model as well as the deviation between reconstructed images of encoder and decoder. We show how much each decoder network component contributes to the reconstructed image deviation.
@@
Table 3: BD-rate (%) results compared to the of floating point model anchor when activations are clipped to fixed range.
@@ 5 DISCUSSION
In this paper, with the aim to achieve deterministic output pictures at encoder and decoder at cross platform tests, we present a method which quantizes the decoder subnetworks of a learned image codec JPEG AI base operating point. We study a simple post training quantization with per layer activation quantization and per channel weights quantization which eliminates the overflow. With this approach, we can keep the original floating point parameters of the encoder-only subnetworks of the codec.

### 11 Jan 2024 v2

@@ 1 INTRODUCTION
Our aim is to achieve bit-exact reconstructed images at decoder when running on different platforms. For this it is sufficient to only quantize the neural network components at the decoder. For many of the use cases of image coding, the decoder of the image codec is fixed, while the encoder design is flexible. Therefore, we quantize only the decoder in order to provide bit exact reconstruction and keep the encoder side free for complex hardware implementation. We reuse the idea in [16] for post-training quantization of weights and activations without overflow and extend it to all decoder subnetworks of the base operation point of JPEG AI VM (verification model) [17] as shown in Fig. 1. We show that by quantizing only decoder networks by reusing the method to eliminate the overflow of the 32-bit accumulator for 16-bit weights and 16-bit activations, we could achieve the deterministic decoded reconstructed images with a minor loss. Also, we provide more experiments to study the 8-bit weights with 16-bit activations, and also to compare the effect of choosing the optimal clipping range instead of clipping to the maximal range for both cases with 8-bit and 16-bit weights. In Section 2, we summarize the previous related work and compare it to ours. The overall description of the learned image coding architecture is given in Section 3.1, and its quantized components are explained in Section 3.2. In Section 3.3, the method of quantization is described. The experimental results are provided in Section 4 and conclusions are given in Section 5.
Fig. 1: High-level architecture diagram of the JPEG AI base operating point. Encoder side subnetworks are $g_{a}$, $h_{a}$, $h_{\mu}$, $h_{\sigma}$. Decoder side subnetworks are $g_{s}$, $h_{\mu}$, $h_{\sigma}$.
@@ 2 RELATED WORK
In [14, 15, 16] we see the quantization of weights and activations in entropy part as a solution to the visible cross platform decoding error. In this paper we extend the post training quantization method of 16-bit weights and activations with no-overflow of 32-bit accumulator in [16]. While [16] quantizes the entropy part only, this work applies quantization to the whole decoder. We target a zero deviation and achieve exact reconstructed images at cross platform tests while keeping the performance loss compared to the floating point model minimal. Although the quantization of only the entropy part as in [16] eliminates the broken images error at the decoder, there still exists a tiny deviation of the quality of the reconstructed images due to the remaining unquantized layers of the decoder, which is not visible visually but can be measured numerically. This is not a critical problem for the image compression task, but for some error-critical applications such as medical image processing, or the video compression task, where the reconstructed image typically becomes a reference for the prediction of the other frames, it is desired to have no deviation. Also we show the results of quantizing only the hyper decoder networks. This allows exact decoding of $\hat{y}$ and can be interesting for computer vision tasks which can use decoded latent tensors $\hat{y}$ directly as input, but do not need the reconstructed image.
In [10, 11, 12, 15], both weights and activations of the whole codec are quantized. In [13] only weights (not activations) of the whole codec are quantized. In [14] weights of the whole codec and activations of only the entropy part are quantized. We quantize weights and activations of whole decoder, i.e., decoder layers in main and entropy part. It is not necessary to quantize all layers in the encoder to achieve the deterministic reconstructed image. This way we allow more complexity at the calculations but the problem of the bit-exact reconstruction is still simple to solve by a post-training quantization method. In common use cases such as streaming, encoding is done once on a device, and decoding is done on many devices. With proposed decoder quantization, for the same input image, encoders on different devices can generate bitstreams with small variation. However, we ensure that given the same bitstream, decoders on different devices generate bit-exact reconstructed images.
@@ 3.1 Learned image codec
As a baseline image codec, we use the base operating point of JPEG AI VM software version 3.4.0 [17]. The JPEG AI VM has two different operating points, high and base, with a different trade-off between complexity and performance. The base operating point architecture consists of an entropy model without context and it has simplified layers and activation functions compared to the high operating point. The architecture of the base operating point for one colour component is shown in Fig. 1. This codec processes colour components separately. Each subnetwork consists of two separate components of subnetworks, one processing the luma channel (Y) and one processing the two chroma channels (U, V). The decoder consists of three main subnetworks, a hyper scale decoder, $h_{\sigma}$, a hyper decoder, $h_{\mu}$, and a main decoder, $g_{s}$; the subnetworks are highlighted in green in Fig. 1. There are four models which are trained with different rate-distortion (RD) loss and overall they produce five different quality points during inference. G-unit and IG-unit are the trained vectors for gain and inverse gain unit, which provide the variable rate functionality as in [18] and thus derive the highest (fifth) rate point from the model of the fourth rate point. The JPEG AI VM software has additional tools to improve the compression quality but all these additional tools are turned off during our implementation, so the coding pipeline is reduced to the neural network layers which are shown in Fig. 2. The explicit block diagram of the light-weighted residual activation unit (ResAU) and the light-weighted residual block (ResBlock) are shown on the right-hand-side of Fig. 2.
@@ 3.2 Quantized part of the codec
We use the floating point trained models as the anchor, and apply the quantization to the decoder subnetworks. The layers which are highlighted in green in Fig. 2 are the convolutional layers to be quantized. The $h_{\sigma}$ is the subnetwork which is calculating the entropy parameters and is shared between the encoder and decoder. From the previous work [15, 14, 16], we know that it is necessary to quantize the entropy part of the subnetwork in order to eliminate catastrophic decoded images. In this paper, we extend the partial quantization of the codec to all the elements of the decoder, with the aim of deterministic calculation of the reconstructed image $\hat{x}$ at the encoder and decoder side. We replace the LeakyReLU activations in the hyper decoder $h_{\mu}$ with an alternative implementation using bit shift. The subnetworks which only exist at the encoder are the main encoder $g_{a}$ and the hyper encoder $h_{a}$; these subnetworks are kept as floating point in our study.
@@ 3.3 Methodology for quantization of decoder subnetworks weights and activations
In [16] weights and activations are both 16-bit. There is one difference of the weights quantization method from [16] if we apply 8-bit quantization to weights. Instead of directly scaling the weights with the calculated factor that eliminates the overflow $2^{-k_{j}}$ as in [16], we use this as an upper bound and search the power of two of scaling factor in interval $[0,k_{j}]$ that results in the minimal quantization error for the weights. We observe that this is necessary to minimize the clipping error. Because we clip the 8-bit weights to the interval $[-2^{7},2^{7}-1]$ instead of $[-2^{15},2^{15}-1]$ as we do for the 16-bit weights case, scaling with the maximum available scaling factor is not ideal for the smaller clipping range.
@@ 4 EXPERIMENTAL RESULTS
The trained models of the JPEG AI VM base operation point are the anchor of the experiments. Then, we apply quantization to the weights and activations of the subnetworks of these models. We use the training, validation and test data which are provided by the JPEG AI common test conditions [20, 21]. For quantization we use the calibration set as in [17]. It is a set of 16 images of various resolutions and content picked from validation set. For the results evaluation, we run the inference on the 50-image JPEG AI test set. We measure the average BD-rate deviation from the floating point model as well as the deviation between reconstructed images of encoder and decoder. We show how much each decoder network component contributes to the reconstructed image deviation.
@@
Table 3: BD-rate (%) results compared to the of floating point model anchor when activations are clipped to fixed range
@@ 5 DISCUSSION
In this paper, with the aim to achieve deterministic output pictures at encoder and decoder at cross platform tests, we present a method which quantizes the decoder subnetworks of a learned image codec JPEG AI base operating point. We study a simple post training quantization with per layer activation quantization and per channel weights quantization which eliminates the overflow. With this approach, we can keep the original floating point parameters of the encoder-only subnetworks of the codec, thereby not limit encoder RD optimization algorithms.